OpenStudy (anonymous):

How many 3 digit numbers have no repeated digits?

OpenStudy (anonymous):

wrong. :(

OpenStudy (angela210793):

i know...sorry

OpenStudy (anonymous):

I think the easiest way would be to figure out how many do have repeated digits, and then take that number away from 900. So look for all the numbers of the form: XXY XYX YXX For example, an XXY number would be something like 336, or 771. XYX would be like 595, or 424. YXX could be like 766, or 388

OpenStudy (angela210793):

ok u have 100 101 110 111 112 113 114 115 116 117 118 119 121 122 131 133 141 144 151 155 161 166 171 177 181 188 191 199 200 202 211 212 ............................................................. idk can't find any f(x) 2 express tht

OpenStudy (anonymous):

Oh, also count all the numbers that are XXX.

OpenStudy (anonymous):

the answer is 648

OpenStudy (angela210793):

till 999 actually

OpenStudy (angela210793):

how did u get tht?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

Til 999, but there are 900 numbers between 100 and 999. I got 647, so I guess I made a calculation error somewhere.

OpenStudy (anonymous):

Yep, found my error. I wrote that there were 10 numbers of the form XXX, when there are in fact 9. Silly me! Anywho, did you give it a shot mynameisanon?

OpenStudy (anonymous):

This is actually a problem related to permutations. First you consider the number of ways of arranging 10 objects (0-9) in three positions. We have: \(10\cdot 9\cdot 8\) Then you subtract the number three digit permutations with zeros first We have: \(9\cdot 8\) The final total is: \(10\cdot 9 \cdot 8 - 9 \cdot 8 = 648\)

OpenStudy (anonymous):

I really think this is what they intended you to do instead of manually counting.

OpenStudy (anonymous):

In any case let me know if you have any questions.